Sanweeve:

A lawn-mowing company is trying to grow its business. It had 30 clients when they started its business and wants to increase by 6 new clients each week. Use an arithmetic sequence to write a function to represent this real-world situation and determine the range of the function for the first four weeks of data. f(x) = 6x + 30; 0 ≤ y ≤ 4 f(x) = 6x + 24; 0 ≤ y ≤ 4 f(x) = 6x + 30; 30 ≤ y ≤ 48 f(x) = 6x + 24; 30 ≤ y ≤ 48

Sanweeve:

what do you mean

darkknight:

So basically, they start with 30 clients, so 30 is the y-intercept. Now it wants to increase by 6 new clients per week so that is the slope Also we want to determine the range from week zero to week 4. Given these 3 pieces of info what can you say?

darkknight:

also one more thing f(x) = 6x + 30; 0 ≤ y ≤ 4 f(x) = 6x + 24; 0 ≤ y ≤ 4 f(x) = 6x + 30; 30 ≤ y ≤ 48 f(x) = 6x + 24; 30 ≤ y ≤ 48 For this part above, is it a typo or does it actually say 0 < y < 4 not x instead of y?

Sanweeve:

i think it would be the second answer

darkknight:

hmm, why do you say that?

darkknight:

Since we need to find the range for the first 4 weeks, it is not answer choice 2, this is because we want to find the range (also this is the most wrong answer choice, if you payed attention it doesn't have the correct y-intercept), if we plug in 0 for x we get 30, if we plug 4 we get 54. Now idk why 30 < y < 54 isn't an answer choice so I am assuming that you typed the thing wrong? as in instead of 0< y < 4 you meant x, (not y) so in that case what is the correct answer choice?

darkknight:

I can't further help if your gonna leave... Just saying ¯\_(ツ)_/¯